When a rigid body rotates about a fixed axis, all the points in the body have the same
1 : Tangential speed.
2 : angular acceleration.
3 : tangential acceleration.
4 : linear displacement.
5 : centripetal acceleration.
1
Expert's answer
2020-05-28T13:09:43-0400

Correct answer is option 2.Angular acceleration.

Reason:

Since,angular acceleration

"\alpha=\frac{d\omega}{dt}=\frac{d^2\theta}{dt^2}"

Thus, all the points on the body traces same angle(angular displacement) per unit time per unit time, which does not depends on how far are those points from axis of rotation.

But,for other options, those quantities explicitly depends on how the points are distributed on the body.